# MSDformer: Multi-scale Discrete Transformer For Time Series Generation

### v1

@@ II Experimental Settings
$[512,512]$ 0.010±.011 0.243±.003 0.032±.002
TABLE XI: The detailed architecture of the time serie tokenizer’s encoder. The ellipsis indicates that layers 3 to 5 are repeated, with the number of repetitions increasing by one for each doubling of the downsampling factor of the encoder.
@@
1 Convolution input channel=$d$, output channel=D, kernel size=3, stride=1, padding=1
2 ReLU nn.ReLU()
3 Convolution input channel=D, output channel=D, kernel size=4, stride=2, padding=1
4 ResNet input channel=D, depth=3, dilation growth rate=3
5 ReLU nn.ReLU()
N Convolution input channel= $d_{c}$, output channel=H, kernel size=3, stride=1, padding=1
TABLE XII: The detailed architecture of the time serie tokenizer’s decoder. The ellipsis indicates that layers 3 to 5 are repeated, with the number of repetitions increasing by one for each doubling of the downsampling factor of the encoder.
Layer Function Descriptions
1 Convolution input channel=$d_{c}$, output channel=D, kernel size=3, stride=1, padding=1
2 ReLU nn.ReLU()
3 ResNet input channel=D, depth=3, dilation growth rate=3
4 ReLU nn.ReLU()
5 Upsample nn.Upsample()
6 Convolution input channel=D, output channel=D, kernel size=3, stride=1, padding=1
N-2 Convolution input channel=D, output channel=D, kernel size=3, stride=1, padding=1
N-1 ReLU nn.ReLU()
N Convolution input channel=D, output channel=$d$, kernel size=3, stride=1, padding=1
III Baselines
@@
IV Additional experimental results
@@
TABLE XIII: The detailed architecture of the Autoregressive Transformer block.
Layer Function Descriptions
1 Layernorm nn.LayerNorm()
2 Casual-attention CasualAttention(q=x, k=x, v=x)
3 Layernorm nn.LayerNorm()
4 MLP nn.Linear()
5 ReLU nn.ReLU()
6 MLP nn.Linear()
Algorithm 2 Training of multi-scale time series tokenizer.
Input: Time series dataset $D=\{x_{1:\tau}^{i}\}_{i=1}^{n}$
1:for $i\leftarrow 1$ to $N_{iter}$ do
2:     Get the $X_{1:\tau}$ $\sim$ $D$;
3:     Initialize residual input $f_{1:\tau}^{(1)}\leftarrow X_{1:\tau}$;
4:     for $k\leftarrow 1$ to $K$ do
5:         Get the $h_{1:L^{(k)}}^{(k)}$ by Equation (2);
6:         Get the $y_{1:L^{(k)}}^{(k)}$ by Equation (3);
7:         Get the $\tilde{h}_{1:L^{(k)}}^{(k)}$ by Equation (4);
8:         Get the $\tilde{f}_{1:\tau}^{(k)}$ by Equation (5);
9:         Get the $f_{1:\tau}^{(k+1)}$ by Equation (6);
10:     end for
11:     Get the $\tilde{x}_{1:\tau}$ by Equation (7);
12:     Compute the training loss $\mathcal{L}$ by Equation (8);
14:     Update the codebook by Equations (9) and (11);
15:end for
1:for $i\leftarrow 1$ to $N_{iter}$ do
3:     Get the label $y_{1:L}$ by Equations (12) and (13);
4:     Get the input $\tilde{y}_{1:L}^{in}$ by Equations (14) and (15);
5:     Feed the $\tilde{y}_{1:L}^{in}$ to $\mathcal{G}_{ms}$;
6:     Compute the training loss $\mathcal{L}_{ms}$ by Equation (22);
8:end for
9:Return trained $\mathcal{G}_{ms}$.
V Algorithms
In this section, we detail the training algorithms for MSDformer, which consists of two stages. In stage 1, our objective is to train a multi-scale time series tokenizer that can produce high-quality discrete representations of time series data while ensuring excellent reconstruction performance. In stage 2, we aim to train a multi-scale autoregressive discrete Transformer that learns the distribution of multi-scale discrete token sequences, thereby enabling the generation of new time series data. Specifically, Algorithm 2 outlines the training procedure for the multi-scale time series tokenizer in stage 1, while Algorithm 3 elucidates the training processes for the multi-scale autoregressive discrete Transformer in stage 2. The inference algorithm is presented in Algorithm 1 within the main text. For the training algorithms for SDformer, simply set the number of scales $K$ in MSDformer to 1.
VI Model details
@@
In this section, we present the detailed network architecture of SDformer and MSDformer. In Stage 1, MSDformer and SDformer employ identical encoder architectures and identical decoder architectures. The only structural variation arises from differing layer counts in specific modules due to distinct downsampling factors, as documented in Tables XI and XII. In stage 2, the multi-scale autoregressive Discrete Transformer is implemented as a standard decoder-only Transformer, with its block structure depicted in Table XIII.

### v2

@@ Appendix B Experimental Settings
$[512,512]$ 0.010±.009 0.243±.003 0.032±.002
TABLE XI: The detailed architecture of the Autoregressive Transformer block.
@@
1 Layernorm $n$.LayerNorm()
2 Casual-attention CasualAttention(q=x, k=x, v=x)
3 Layernorm $n$.LayerNorm()
4 MLP $n$.Linear()
5 ReLU $n$.ReLU()
6 MLP $n$.Linear()
TABLE XII: The detailed architecture of the time serie tokenizer’s encoder. The ellipsis indicates that layers 3 to 5 are repeated, with the number of repetitions increasing by one for each doubling of the downsampling factor of the encoder.
Layer Function Descriptions Input Shape Output Shape
1 Convolution input channel=$d$, output channel=D, kernel size=3, stride=1, padding=1 (B, $d$, L) (B, D, L)
2 ReLU $n$.ReLU() (B, D, L) (B, D, L)
3 Convolution input channel=D, output channel=D, kernel size=4, stride=2, padding=1 (B, D, L) (B, D, L//2)
4 ResNet input channel=D, depth=3, dilation growth rate=3 (B, D, L//2) (B, D, L//2)
$\dots$ $\dots$ Repeat layers 3-4 for (T-1) more times (B, D, $\text{L}/2^{i}$) (B, D, $\text{L}/2^{i+1}$)
[…]
TABLE XIII: The detailed architecture of the time serie tokenizer’s decoder. The ellipsis indicates that layers 3 to 5 are repeated, with the number of repetitions increasing by one for each doubling of the downsampling factor of the encoder.
Layer Function Descriptions Input Shape Output Shape
[…]
2 ReLU $n$.ReLU() (B, D, $\text{L}/2^{T}$) (B, D, $\text{L}/2^{T}$)
3 ResNet input channel=D, depth=3, dilation growth rate=3 (B, D, $\text{L}/2^{T-i}$) (B, D, $\text{L}/2^{T-i}$)
4 Upsample nn.Upsample(scale factor=2, mode=’nearest’) (B, D, $\text{L}/2^{T-i}$) (B, D, $\text{L}/2^{T-i-1}$)
5 Convolution input channel=D, output channel=D, kernel size=3, stride=1, padding=1 (B, D, $\text{L}/2^{T-i-1}$) (B, D, $\text{L}/2^{T-i-1}$)
Repeat layers 3-5 for (T-1) more times
3T+3 Convolution input channel=D, output channel=D, kernel size=3, stride=1, padding=1 (B, D, L) (B, D, L)
3T+4 ReLU $n$.ReLU() (B, D, L) (B, D, L)
3T+5 Convolution input channel=D, output channel=$d$, kernel size=3, stride=1, padding=1 (B, D, L) (B, $d$, L)
TABLE XIV: The detailed architecture of the ResNet block.
Layer Function Descriptions Input Shape Output Shape
1 LayerNorm $n$.LayerNorm() (B, D, L) (B, D, L)
2 ReLU $n$.ReLU() (B, D, L) (B, D, L)
3 Convolution input channel=D, output channel=D, kernel size=3, padding=dilation, dilation=dilation (B, D, L) (B, D, L)
4 LayerNorm $n$.LayerNorm() (B, D, L) (B, D, L)
5 ReLU $n$.ReLU() (B, D, L) (B, D, L)
6 Convolution input channel=D, output channel=D, kernel size=1, stride=1, padding=0 (B, D, L) (B, D, L)
7 Residual output = input + conv output (B, D, L) (B, D, L)
TABLE XV: The detailed training process of stage 2.
Process Module Operation Input Shape Output Shape
Input Codebook indices inputs (B, $\text{P}_{\text{sum}}$)
[…]
Appendix C Baselines
@@
Appendix D Additional experimental results
@@
Appendix E Algorithms
In this section, we summarize the complete training and inference procedures of MSDformer, which follows a two-stage pipeline and are shown in the main manuscript. Stage 1 trains a multi-scale time-series tokenizer to discretize continuous sequences into compact code indices while preserving reconstruction fidelity, providing the discrete representation on which the rest of the framework is built. Stage 2 trains a multi-scale autoregressive Transformer to model the joint distribution of the resulting multi-scale token sequences, which enables unconditional sampling of new tokens and, through the tokenizer’s decoder, the generation of novel time-series samples. The detailed optimization steps for Stage 1 and Stage 2 are summarized in Algorithms 1 and 2, respectively, while the overall inference pipeline is given in Algorithm 3. SDformer is a special case of this framework obtained by setting the number of scales to (KK=1).
Appendix F Model details
@@
In this section, we present the detailed network architecture of SDformer and MSDformer.The meanings of the symbols in following tables are as follows: B is batch size, $d$ is the input channels, D and $d_{c}$ denote the hidden and codebook dimensions, respectively. L, T are the sequence length and number of downsampling/upsampling operations in our setting. Also the i in Table XII and XIII represents iteration index (0 to T-1). In Stage 1, MSDformer and SDformer employ identical encoder architectures and identical decoder architectures. The only structural variation arises from differing layer counts in specific modules due to distinct downsampling factors, as documented in Tables XII and XIII. Also, the ResNet block of Stage 1 is described in Table XIV. In stage 2, the multi-scale autoregressive Discrete Transformer is implemented as a standard decoder-only Transformer, with its block structure depicted in Table XI and the detailed training process is shown in Table XV, in which the $\text{P}_{\text{sum}}$ is sum of the token length within each scale and $\text{C}_{\text{sum}}$ denotes the total codebook size of the multi-scale tokenizers. Since the T denotes the number of upsampling and downsampling operations. The total layers of Encoder include 2 initial layers (Conv + ReLU), 2*T repeated layers (Conv + ResNet) and 1 final layer (Conv). For the decoder structure, it consists of 2 initial layers (Conv + ReLU), 3*T repeated layers (ResNet + Upsample + Conv) and 3 final layers (Conv + ReLU + Conv).
Finally, the multi-scale codebook sizes $V^{(1:K)}$ used by MSDformer are selected empirically per dataset. Since different datasets exhibit different dimensionality, which requires different effective token rate budgets for a favorable rate–distortion trade-off. As future work, we plan to explore the relationship between the optimal token budget (e.g., effective rate ($\sum_{k}L^{(k)}\log_{2}V^{(k)})$) and dataset dimensionality, with the goal of deriving an automatic and theoretically grounded configuration strategy for $V^{(1:K)}$.
MSDformer
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
((a))
((b))
((c))
((d))
((e))
((f))
SDformer
Diffusion-TS
MSDformer
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
((a))
((b))
((c))
((d))
((e))
((f))
SDformer
Diffusion-TS